• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 16:05:32 on April12,2007

    Entry number 198758

    keyword=Shift Summary

    Started shift with the hall in controlled access while the right Q1
    was fixed. By 11:30 the problem was solved and beam studies which needed the hall began. By 14:30 over 100 uA of beam could be sent to hall A and beam was given back to our control.

    Start/Stop Run Target Notes
    14:30/15:39 2368 Al Dummy 40 uA
    15:57/NextShift 2369 3He 100 uA